# Quasi-boolean groups

### v2

@@ 3. $(iii)$
Clearly $G$ is the maximal pro-$2$ quotient of $\overline{G}$. On the other hand, $\overline{G}$ is a real free profinite group. In fact, $\overline{G}$ is the group $\widehat{D}(X,Y_{+},e)$ in [6, Lemma 1.3 on pages 156–157], where $Y_{+}$ is the one-point compactification of $Y$ equipped with the discrete topology, and $e\in Y_{+}-Y$ is the point at infinity. This finishes the proof of Theorem 6.6. ∎
@@
The following lemma is probably very well-known, but we could not find a convenient reference:
@@ 2. $(ii)$
The Pontryagin dual of $G$ is a discrete $\mathbb{F}_{p}$-linear vector space $V$, since $G$ is compact. Note that every isomorphism between discrete $\mathbb{F}_{p}$-linear vector spaces is automatically a homeomorphism, so $V$ is isomorphic to the direct sum $\mathbb{F}_{p}^{\oplus X}$, equipped with the discrete topology, as a topological group for some set $X$. The Pontryagin dual of $\mathbb{F}_{p}^{\oplus X}$ is $\mathbb{F}_{p}^{X}$, which is isomorphic to $G$ by Pontryagin duality. So claim $(i)$ holds.
Assertion $(ii)$ is equivalent to the following claim: let $i\colon B\to A$ be a monomorphism of $p$-torsion abelian profinite groups. Then there is a morphism $j\colon A\to B$ such that $j\circ i=\mathrm{id}_{B}$. By Pontryagin duality, it is equivalent to the following claim: let $p\colon V\to U$ be an epimorphism of $\mathbb{F}_{p}$-linear vector spaces. Then there is a morphism $r\colon U\to V$ such that $p\circ r=\mathrm{id}_{V}$. The latter is well-known. ∎
Again let $G$ be a quasi-Boolean pro-$2$ group which is the free product of a free pro-$2$ group $F$ and a Boolean group $\mathbb{B}(X)$ for a profinite space $X$. Let $j_{X}\colon X\to\mathcal{X}(\mathbb{B}(X))$ denote the composition of the map $i_{X}\colon X\to\mathcal{Y}(\mathbb{B}(X))$, the inclusion $\mathcal{Y}(\mathbb{B}(X))\subseteq\mathcal{Y}(G)$, and the quotient map $\mathcal{Y}(G)\to\mathcal{X}(G)$. It is useful to record the following fact:
@@
The map $j_{X}$ is a homeomorphism onto $\mathcal{X}^{*}(G)$.
@@
Since $X$ is compact and $\mathcal{X}^{*}(G)$ is Hausdorff, it will be sufficient to show that $j_{X}$ maps $X$ onto $\mathcal{X}^{*}(G)$ bijectively. Let $x,y\in X$ be two different elements. Then there is a continuous map $s\colon X\to\mathbb{Z}/2$ such that $s(x)\neq s(y)$. Let $\overline{s}\colon G\to\mathbb{Z}/2$ be the unique homomorphism such that the restriction of $\overline{s}$ onto $F$ is trivial, and onto $\mathbb{B}(X)$ is $b_{s}$ (as defined in Remark 6.3). Under $\overline{s}$ the images of $i_{X}(x)$ and $i_{X}(y)$ are not conjugate, so they are not conjugate in $G$ either. Therefore $j_{X}$ is injective. Since we may guarantee that $s(x)$ is not a unit, we get that $j_{X}$ maps into $\mathcal{X}^{*}(G)$, too.
Now assume that there is a $y\in\mathcal{X}^{*}(G)$ which is not in the image of $j_{X}$. Since $X$ is compact and $\mathcal{X}^{*}(G)$ is Hausdorff, the image of $j_{X}$ is closed, so by Lemma 3.6 there is a continuous function $r\colon\mathcal{X}^{*}(G)\to\mathbb{Z}/2$ such that $r\circ j_{X}$ is zero, and $r(y)$ is non-zero. By Theorem 6.6, the pro-$2$ group $G$ is real projective. Hence, by Scheiderer’s theorem [8, Theorem 2.11], there is a continuous homomorphism $\overline{r}\colon G\to\mathbb{Z}/2$ whose image under the map
$\pi_{1}\colon H^{1}(G,\mathbb{Z}/2)\to C(\mathcal{X}^{*}(G),\mathbb{Z}/2)$
is $r$. Then the restriction of $\overline{r}$ onto $i_{X}(X)$ is zero, so by the universal property of $\mathbb{B}(X)$ the restriction of $\overline{r}$ onto $\mathbb{B}(X)$ is also zero, and hence this homomorphism factors through the surjective homomorphism $p_{1}\colon G\to F$ supplied by the universal property of free pro-$2$ products. But $F$ is torsion-free (see [12, Corollary 3, part (a) on page 31]), so $p_{1}$ is zero on any involution $\overline{y}$ in the conjugacy class $y$. Therefore $r(y)=\overline{r}(\overline{y})$ is zero, which is a contradiction. ∎
@@
Let $H\subseteq G$ be the open subgroup corresponding to the finite extension $K(\sqrt{-1})/K$. Then $H$ is isomorphic to the absolute Galois group of $K(\sqrt{-1})$, and it is torsion-free. In particular, if $H=G$ then $\mathcal{X}^{*}(G)$ is empty and the claim is trivially true. Otherwise, $H$ has index two in $G$. Then the theorem follows from Proposition 6.13 below and Theorem 4.7. ∎
@@
Clearly $\mathcal{Y}(G)$ is closed in $G$. Since $H$ is torsion-free, the subset $\mathcal{Y}^{*}(G)$ is the intersection of $\mathcal{X}(G)$ and the complement of the open $H$ in $G$, so it is closed, too. Since $G$ is profinite, we get that $\mathcal{Y}^{*}(G)$ is profinite. As $G$ is isomorphic to an absolute Galois group, for every $y\in\mathcal{Y}^{*}(G)$ we have $H\cap C_{G}(y)=\{1\}$, and hence $H$ acts freely on $\mathcal{Y}^{*}(G)$. This action is also clearly continuous.
We claim that every $x\in G$ conjugate to $y$ is already conjugate under $H$. Indeed, let $z\in G$ be such that $z^{-1}xz=y$. Since $H$ has index $2$ and $y\not\in H$ we have $G=H\cup Hy$. If $z\in H$ the claim is clearly true. Otherwise $z=hy$ for some $h\in H$, and hence $x=zyz^{-1}=hyyy^{-1}h^{-1}=hyh^{-1}$, so $x$ is conjugate to $y$ under $H$ in this case, too. So the map $\mathcal{Y}^{*}(G)\to\mathcal{X}^{*}(G)$ is the quotient map with respect to the action of $H$. Since $\mathcal{Y}^{*}(G)$ is open in $\mathcal{Y}(G)$ the subspace topology on $\mathcal{X}^{*}(G)\subset\mathcal{X}(G)$ is the quotient topology with respect to the map in the claim. Therefore, $\mathcal{X}^{*}(G)$ is Hausdorff by Proposition 4.5. ∎
@@
Since every real projective profinite group is isomorphic to the absolute Galois group of a pseudo real closed field, the claim follows at once from Theorem 6.12. ∎
@@
Note that Corollary 6.14 is also part $(a)$ of [6, Lemma 3.5 on page 160]. We think, however, that our proof is more conceptual and derives a similar claim for a much larger class of groups.
Notation 6.16.
@@
Proof.
Let $X$ denote $\mathcal{X}^{*}(G)$, and let $s\colon X\to\mathcal{Y}^{*}(G)$ be the section furnished by Theorem 6.12. Let $B$ be the image of the homomorphism $(b_{s})_{*}\colon\mathbb{B}(X)_{*}\to G_{*}$. Since $\mathbb{B}(X)_{*}$ is compact and $G_{*}$ is Hausdorff, $B$ is closed. Let $A\subseteq G_{*}$ be a complement of $B$, which exists by part $(ii)$ of Lemma 6.10. By part $(i)$ of Lemma 6.10, there is a set $Y$ such that $A\cong\mathbb{Z}/2^{Y}$. Since $F(Z)_{*}\cong\mathbb{Z}/2^{Y}$, by part (a) of [12, Proposition 24 on page 30] and Pontryagin duality, there is a continuous homomorphism $h\colon F(Y)\to G$ such that $h_{*}\colon F(Y)_{*}\to G_{*}$ maps $F(Y)_{*}$ isomorphically onto $A$. We set $P=F(Y)*_{2}\mathbb{B}(X)$ and let $\alpha\colon P\to G$ be the homomorphism $h*_{2}b_{s}$. Since $(G_{1}*_{2}G_{2})_{*}\cong(G_{1})_{*}\oplus(G_{2})_{*}$ for every pair of pro-$2$ groups $G_{1}$ and $G_{2}$, we get that $\alpha_{*}\colon P_{*}\to G_{*}$ is an isomorphism. Hence, by [12, Proposition 24, part (b), on page 30] and Pontryagin duality, the map $\alpha$ is surjective. To finish the proof we need the following
@@
This claim is part (b) of [6, Lemma 3.5 on page 160]. The proof relies on the projectivity of the Artin–Schreier structures attached to real projective groups (see [5, Proposition 7.7 on page 473]). ∎
We return to the proof of Theorem 6.17. By Theorem 6.6, the quasi-Boolean group $P$ is real projective, while by Theorem 6.11 the subset $X\subseteq P$ is a system of representatives of $\mathcal{X}^{*}(P)$ mapped bijectively onto a system of representatives of $\mathcal{X}^{*}(G)$, so the conditions in Lemma 6.18 above hold. Therefore there is a continuous homomorphism $\gamma\colon G\to P$ such that $\alpha\circ\gamma=\mathrm{id}_{G}$. Then $\alpha_{*}\circ\gamma_{*}=(\alpha\circ\gamma)_{*}=(\mathrm{id}_{G})_{*}=\mathrm{id}_{G_{*}}$ is the identity, and $\alpha_{*}$ is an isomorphism, so $\gamma_{*}$ is an isomorphism, too. Therefore, $\gamma$ is surjective by part (b) of [12, Proposition 24 on page 30] and Pontryagin duality. So $\gamma$ is an isomorphism, and hence $G$ is quasi-Boolean. This finishes the proof of Theorem 6.17. ∎
@@ 7. Quillen’s theorems and their consequences
Following Quillen we say that a homomorphism $R\to S$ of graded anti-commutative rings is finite if $S$ is a finitely generated module over $R$. This definition is a bit ambiguous since it does not specify whether we consider $S$ as a left $R$-module or a right $R$-module. However, note that $S$ is a finitely generated left $R$-module if and only if it is a finitely generated right $R$-module. Indeed if $S$ is a finitely generated left $R$-module then it is also generated by a finite set $H\subset S$ of homogeneous elements. But the left $R$-module generated by $H$ is the same as the right $R$-module generated by $H$ since $S$ is anti-commutative. Therefore $S$ is also finitely generated as a right $R$-module. The converse could be proved similarly.
@@
Let $G$ be a pro-$p$ group and let $H\subseteq G$ be a finite subgroup. Then the homomorphism $H^{\bullet}(G,\mathbb{Z}/p)\to H^{\bullet}(H,\mathbb{Z}/p)$ is finite.
@@
Note that $H$ is a closed subgroup, as $G$ is Hausdorff, so the homomorphism $H^{\bullet}(G,\mathbb{Z}/p)\to H^{\bullet}(H,\mathbb{Z}/p)$ is well-defined. This theorem was proved by Quillen when $G$ is finite [13, Corollary 2.4 on page 555], and the general case follows as an easy corollary. Indeed let $N\triangleleft G$ be an open normal subgroup such that the restriction of the quotient map $q\colon G\to G/N$ to $H$ is injective. Then the homomorphism $H^{\bullet}(G/N,\mathbb{Z}/p)\to H^{\bullet}(H,\mathbb{Z}/p)$ induced by the composition of the inclusion map $H\to G$ and $q$ is finite by the above. Since this homomorphism factors through $H^{\bullet}(G,\mathbb{Z}/p)\to H^{\bullet}(H,\mathbb{Z}/p)$, the latter is also finite. ∎
@@
Let $G$ be a pro-$p$ group and let $H\subseteq G$ be a subgroup of order $p$. Then the homomorphism $H^{n}(G,\mathbb{Z}/p)\to H^{n}(H,\mathbb{Z}/p)$ is non-zero for infinitely many $n$.
@@
Assume that the claim is false and there is a natural number $d$ such that the image of $H^{n}(G,\mathbb{Z}/p)\to H^{n}(H,\mathbb{Z}/p)$ is zero for $n\geq d$. Let $S\subset H^{\bullet}(H,\mathbb{Z}/p)$ be a finite subset of homogeneous elements which generate $H^{\bullet}(H,\mathbb{Z}/p)$ as a $H^{\bullet}(G,\mathbb{Z}/p)$-module. Let $d^{\prime}$ be the maximal degree of the elements of $S$. Then $H^{n}(H,\mathbb{Z}/p)=0$ for every $n\geq d+d^{\prime}$. But, since $H$ has order $p$, $H^{n}(H,\mathbb{Z}/p)\neq 0$ for every $n$ which is a contradiction. ∎
@@
Let $G$ be a cohomologically quasi-Boolean pro-$2$ group. Then the quotient $H^{\bullet}(G,\mathbb{Z}/2)/\mathcal{N}(H^{\bullet}(G,\mathbb{Z}/2))$ is a graded Boolean algebra.
@@
Let $H^{*}(G,\mathbb{Z}/2)$ be the connected sum of a dual algebra $D^{*}$ and a graded Boolean algebra $B^{*}$. Then the nilradical of $H^{\bullet}(G,\mathbb{Z}/2)$ is $D^{1}$, so the quotient $H^{\bullet}(G,\mathbb{Z}/2)/\mathcal{N}(H^{\bullet}(G,\mathbb{Z}/2))$ is isomorphic to $B^{*}$. ∎
@@
Let $G$ be a cohomologically quasi-Boolean pro-$2$ group. Let $B$ denote the up to isomorphism unique Boolean ring such that the associated graded Boolean algebra $B^{*}$ is isomorphic to $H^{\bullet}(G,\mathbb{Z}/2)/\mathcal{N}(H^{\bullet}(G,\mathbb{Z}/2))$. We say that a continuous homomorphism $k\colon G\to\mathbb{Z}/2$ is a quasi-canonical homomorphism if the image of the associated cohomology class $k\in H^{1}(G,\mathbb{Z}/2)$ under the quotient map
$H^{\bullet}(G,\mathbb{Z}/2)\to H^{\bullet}(G,\mathbb{Z}/2)/\mathcal{N}(H^{\bullet}(G,\mathbb{Z}/2))$

### v3

@@ 3. $(iii)$
Clearly $G$ is the maximal pro-$2$ quotient of $\overline{G}$. On the other hand, $\overline{G}$ is a real free profinite group. In fact, $\overline{G}$ is the group $\widehat{D}(X,Y_{+},e)$ in [10, Lemma 1.3 on pages 156–157], where $Y_{+}$ is the one-point compactification of $Y$ equipped with the discrete topology, and $e\in Y_{+}-Y$ is the point at infinity. This finishes the proof of Theorem 6.6. ∎
@@
The following lemma is probably very well-known, but we could not find a convenient reference. For its proof we recall from [14, Definition 1.1.10] that the Pontryagin dual of a Hausdorff, abelian and locally compact group $A$ is the group $A^{\vee}=\mathrm{Hom}_{\mathrm{cont}}(A,\mathbb{R}/\mathbb{Z})$ of continuous group homomorphisms from $A$ to $\mathbb{R}/\mathbb{Z}$ where the latter is equipped with the quotient topology inherited from $\mathbb{R}$. The group $A^{\vee}$ is considered as a topological group with the compact-open topology. We recall that Pontryagin duality states that the canonical homomorphism $A\to(A^{\vee})^{\vee}$ is an isomorphism (see [14, Theorem 1.1.11] and the references therein). In particular, for an abelian profinite group $G$, the Pontryagin dual $G^{\vee}$ is a discrete abelian torsion group.
@@ 2. $(ii)$
The Pontryagin dual of $G$ is a discrete $\mathbb{F}_{p}$-linear vector space $V$ since $G$ is compact. Note that every isomorphism between discrete $\mathbb{F}_{p}$-vector spaces is automatically a homeomorphism, so $V$ is isomorphic to the direct sum $\mathbb{F}_{p}^{\oplus X}$, equipped with the discrete topology, as a topological group for some set $X$. The Pontryagin dual of $\mathbb{F}_{p}^{\oplus X}$ is $\mathbb{F}_{p}^{X}$, which is isomorphic to $G$ by Pontryagin duality. So claim $(i)$ holds.
Assertion $(ii)$ is equivalent to the following claim: let $i\colon B\to A$ be a monomorphism of $p$-torsion abelian profinite groups. Then there is a morphism $j\colon A\to B$ such that $j\circ i=\mathrm{id}_{B}$. By Pontryagin duality, it is equivalent to the following claim: let $p\colon V\to U$ be an epimorphism of $\mathbb{F}_{p}$-linear vector spaces. Then there is a morphism $r\colon U\to V$ such that $p\circ r=\mathrm{id}_{U}$. The latter is well-known. ∎
For later purposes, we record the following fact:
@@
Let $G$ be a quasi-Boolean pro-$2$ group which is the free product of a free pro-$2$ group $F$ and a Boolean group $\mathbb{B}(X)$ for a profinite space $X$. Let $j_{X}\colon X\to\mathcal{X}(\mathbb{B}(X))$ denote the composition of the map $i_{X}\colon X\to\mathcal{Y}(\mathbb{B}(X))$, the inclusion $\mathcal{Y}(\mathbb{B}(X))\subseteq\mathcal{Y}(G)$, and the quotient map $\mathcal{Y}(G)\to\mathcal{X}(G)$. Then the map $j_{X}$ is a homeomorphism onto $\mathcal{X}^{*}(G)$.
To prove the theorem we first show the following
Lemma 6.12.
For a real projective group $G$, the space $\mathcal{X}^{*}(G)$ is compact and totally separated.
@@
Since $G$ is real projective, the subset $\mathcal{Y}^{*}(G)=\mathrm{Inv}(G)$ of involutions is closed in $G$. Hence $\mathrm{Inv}(G)$ is compact and totally separated by $(i)\Rightarrow(iv)$ of Theorem 3.2. This shows that the quotient $\mathcal{X}^{*}(G)$ of conjugacy classes is also compact by [2, I §9.4, Theorem 2 on page 87]. Now let $x,y$ be two distinct involutions in $G$ which are not conjugate. Since $\mathrm{Inv}(G)$ is totally separated, there is an open and closed subset $U\subset\mathrm{Inv}(G)$ such that $x\in U$ and $y\notin U$. By definition of the subspace topology and since the open normal subgroups form a basis for the topology on $G$, there is an open normal subgroup $\widetilde{U}$ of $G$ such that $U=\mathrm{Inv}(G)\cap\widetilde{U}$. Since $\widetilde{U}$ is conjugation-invariant, both $U$ and $\mathrm{Inv}(G)-U$ are conjugation-invariant. This implies that the image of $U$ under the quotient map $\mathrm{Inv}(G)\to\mathcal{X}^{*}(G)$ is open and closed and contains the conjugacy class of $x$ but not the class of $y$. This shows that $\mathcal{X}^{*}(G)$ is totally separated. ∎
Proof of Theorem 6.11.
By Lemma 6.12, $\mathcal{X}^{*}(G)$ is totally separated and hence Hausdorff. Since $X$ is compact, it will therefore be sufficient to show that $j_{X}$ maps $X$ onto $\mathcal{X}^{*}(G)$ bijectively by [2, I §9.4, Corollary 2 on page 87]. Let $x,y\in X$ be two distinct elements. Then there is a continuous map $s\colon X\to\mathbb{Z}/2\mathbb{Z}$ such that $s(x)\neq s(y)$. Let $\overline{s}\colon G\to\mathbb{Z}/2\mathbb{Z}$ be the unique homomorphism such that the restriction of $\overline{s}$ onto $F$ is trivial, and onto $\mathbb{B}(X)$ is $b_{s}$ (as defined in Remark 6.3). Under $\overline{s}$ the images of $i_{X}(x)$ and $i_{X}(y)$ are not conjugate, so they are not conjugate in $G$ either. Therefore $j_{X}$ is injective. Since we may guarantee that $s(x)$ is not a unit, we get that $j_{X}$ maps into $\mathcal{X}^{*}(G)$, too.
Now assume that there is a $y\in\mathcal{X}^{*}(G)$ which is not in the image of $j_{X}$. Since $X$ is compact and $\mathcal{X}^{*}(G)$ is Hausdorff, the image of $j_{X}$ is closed by [2, I §9.4, Corollary 2 on page 87]. Thus, by Lemma 3.7, which applies to $\mathcal{X}^{*}(G)$ by Lemma 6.12, there is a continuous function $r\colon\mathcal{X}^{*}(G)\to\mathbb{Z}/2\mathbb{Z}$ such that $r\circ j_{X}$ is zero, and $r(y)$ is non-zero. By Theorem 6.6, the pro-$2$ group $G$ is real projective. Hence, by Scheiderer’s theorem [15, Theorem 2.11], there is a continuous homomorphism $\overline{r}\colon G\to\mathbb{Z}/2\mathbb{Z}$ whose image under the map
$\pi_{1}\colon H^{1}(G,\mathbb{Z}/2\mathbb{Z})\to C(\mathcal{X}^{*}(G),\mathbb{Z}/2\mathbb{Z})$
is $r$ where $C(\mathcal{X}^{*}(G),\mathbb{Z}/2\mathbb{Z})$ denotes the ring of continuous functions from $\mathcal{X}^{*}(G)$ to $\mathbb{Z}/2\mathbb{Z}$, where the latter is equipped with the discrete topology. Then the restriction of $\overline{r}$ onto $i_{X}(X)$ is zero, so by the universal property of $\mathbb{B}(X)$ the restriction of $\overline{r}$ onto $\mathbb{B}(X)$ is also zero, and hence this homomorphism factors through the surjective homomorphism $p_{1}\colon G\to F$ supplied by the universal property of free pro-$2$ products. But $F$ is torsion-free (see [21, Corollary 4 on page 31]), so $p_{1}$ is zero on any involution $\overline{y}$ in the conjugacy class $y$. Therefore $r(y)=\overline{r}(\overline{y})$ is zero, which is a contradiction. ∎
@@
Let $H\subseteq G$ be the open subgroup corresponding to the finite extension $K(\sqrt{-1})/K$. Then $H$ is isomorphic to the absolute Galois group of $K(\sqrt{-1})$, and hence it is torsion-free by the Artin–Schreier theorem. In particular, if $H=G$ then $\mathcal{X}^{*}(G)$ is empty and the claim is trivially true. Otherwise, $H$ has index two in $G$. Then the theorem follows from Proposition 6.14 below and Theorem 4.7. ∎
@@
Clearly $\mathcal{Y}(G)$ is closed in $G$. Since $H$ is torsion-free, the subset $\mathcal{Y}^{*}(G)$ is the intersection of $\mathcal{Y}(G)$ and the complement of the open $H$ in $G$, so it is closed, too. Let $C_{G}(y)$ denote the centrailzer of $y$ in $G$. Since $G$ is profinite, we get that $\mathcal{Y}^{*}(G)$ is profinite. As $G$ is isomorphic to an absolute Galois group, for every $y\in\mathcal{Y}^{*}(G)$ we have $H\cap C_{G}(y)=\{1\}$, and hence $H$ acts freely on $\mathcal{Y}^{*}(G)$. This action is also clearly continuous.
We claim that every $x\in G$ conjugate to $y$ is already conjugate under $H$. Indeed, let $z\in G$ be such that $z^{-1}xz=y$. Since $H$ has index $2$ and $y\not\in H$, we have $G=H\cup Hy$. If $z\in H$ the claim is clearly true. Otherwise $z=hy$ for some $h\in H$, and hence $x=zyz^{-1}=hyyy^{-1}h^{-1}=hyh^{-1}$, so $x$ is conjugate to $y$ under $H$ in this case, too. So the map $\mathcal{Y}^{*}(G)\to\mathcal{X}^{*}(G)$ is the quotient map with respect to the action of $H$. Since $\mathcal{Y}^{*}(G)$ is open in $\mathcal{Y}(G)$, the subspace topology on $\mathcal{X}^{*}(G)\subset\mathcal{X}(G)$ is the quotient topology with respect to the map in the claim. Therefore, $\mathcal{X}^{*}(G)$ is Hausdorff by Proposition 4.5. ∎
@@
By [9, Theorem 10.4 on page 487], every real projective profinite group is isomorphic to the absolute Galois group of a pseudo real closed field. The claim then follows at once from Theorem 6.13. ∎
@@
Note that Corollary 6.15 is also part $(a)$ of [10, Lemma 3.5 on page 160]. We think, however, that our proof is more conceptual and derives a similar claim for a much larger class of groups.
Notation 6.17.
@@
For the proof of the theorem we recall the following facts from [21, Proposition 24 on page 30] for the special case $p=2$:
Proposition 6.19.
Let $G$ be a pro-$2$ group and $I$ a set. Let
$\theta\colon H^{1}(G,\mathbb{Z}/2\mathbb{Z})\to(\mathbb{Z}/2\mathbb{Z})^{I}$
be a homomorphism. Then:
  • (a)
    There exists a morphism $f\colon F(I)\to G$ such that $\theta=H^{1}(f)$ where $F(I)$ denotes the free pro-$2$ group on $I$.
  • (b)
    If $\theta$ is injective, the morphism $f$ is surjective. ∎
Proof of Theorem 6.18.
Let $X$ denote $\mathcal{X}^{*}(G)$, and let $s\colon X\to\mathcal{Y}^{*}(G)$ be the section furnished by Theorem 6.13. We let $B$ denote the image of the homomorphism $(b_{s})_{*}\colon\mathbb{B}(X)_{*}\to G_{*}$ where $b_{s}\colon\mathbb{B}(X)\to\mathcal{Y}^{*}(G)\subset G$ is the homomorphism defined in Remark 6.3. Since $\mathbb{B}(X)_{*}$ is compact and $G_{*}$ is Hausdorff, $B$ is closed in $G_{*}$ by [2, I §9.4, Corollary 2 on page 87]. Let $A\subseteq G_{*}$ be a complement of $B$, which exists by part $(ii)$ of Lemma 6.10. By part $(i)$ of Lemma 6.10, there is a set $Y$ such that $A\cong(\mathbb{Z}/2\mathbb{Z})^{Y}$. Since $F(Y)_{*}\cong(\mathbb{Z}/2\mathbb{Z})^{Y}$, by part (a) of Proposition 6.19 and Pontryagin duality, there is a continuous homomorphism $h\colon F(Y)\to G$ such that $h_{*}\colon F(Y)_{*}\to G_{*}$ maps $F(Y)_{*}$ isomorphically onto $A$. We set $P=F(Y)*_{2}\mathbb{B}(X)$ and let $\alpha\colon P\to G$ be the homomorphism $h*_{2}b_{s}$. Since $(G_{1}*_{2}G_{2})_{*}\cong(G_{1})_{*}\oplus(G_{2})_{*}$ for every pair of pro-$2$ groups $G_{1}$ and $G_{2}$, we get that $\alpha_{*}\colon P_{*}\to G_{*}$ is an isomorphism. Hence, by part (b) of Proposition 6.19 and Pontryagin duality, the map $\alpha$ is surjective. To finish the proof we need the following
@@
This claim is part (b) of [10, Lemma 3.5 on page 160]. The proof relies on the projectivity of the Artin–Schreier structures attached to real projective groups (see [9, Proposition 7.7 on page 473]). ∎
We return to the proof of Theorem 6.18. By Theorem 6.6, the quasi-Boolean group $P$ is real projective, while by Theorem 6.11 the subset $X\subseteq P$ is a system of representatives of $\mathcal{X}^{*}(P)$ mapped bijectively onto a system of representatives of $\mathcal{X}^{*}(G)$, so the conditions in Lemma 6.20 above hold. Therefore there is a continuous injective homomorphism $\gamma\colon G\to P$ such that $\alpha\circ\gamma=\mathrm{id}_{G}$. Then the composition $\alpha_{*}\circ\gamma_{*}=(\alpha\circ\gamma)_{*}=(\mathrm{id}_{G})_{*}=\mathrm{id}_{G_{*}}$ is the identity, and $\alpha_{*}$ is an isomorphism, so $\gamma_{*}$ is an isomorphism, too. Therefore, $\gamma$ is surjective by part (b) of Proposition 6.19 and Pontryagin duality. So $\gamma$ is an isomorphism, and hence $G$ is quasi-Boolean. This finishes the proof of Theorem 6.18. ∎
@@ 7. Quillen’s theorems and their consequences
Following Quillen we say that a homomorphism $R\to S$ of graded anti-commutative rings is finite if $S$ is a finitely generated module over $R$.
This definition is a bit ambiguous since it does not specify whether we consider $S$ as a left $R$-module or a right $R$-module. However, note that $S$ is a finitely generated left $R$-module if and only if it is a finitely generated right $R$-module. Indeed if $S$ is a finitely generated left $R$-module then it is also generated by a finite set $H\subset S$ of homogeneous elements. But the left $R$-module generated by $H$ is the same as the right $R$-module generated by $H$ since $S$ is anti-commutative. Therefore $S$ is also finitely generated as a right $R$-module. The converse could be proved similarly.
@@
Let $G$ be a pro-$p$ group and let $H\subseteq G$ be a finite subgroup. Then the homomorphism $H^{\bullet}(G,\mathbb{Z}/p\mathbb{Z})\to H^{\bullet}(H,\mathbb{Z}/p\mathbb{Z})$ is finite.
@@
Recall that every finite subset of a Hausdorff space is closed by [2, I §8.2, Proposition 4 on page 77]. Since $G$ is Hausdorff, we conclude that $H$ is a closed subgroup. Hence the homomorphism $H^{\bullet}(G,\mathbb{Z}/p\mathbb{Z})\to H^{\bullet}(H,\mathbb{Z}/p\mathbb{Z})$ is well-defined. The analog of the theorem was proved by Quillen when $G$ is finite [22, Corollary 2.4 on page 555], and we now show that the general case follows as an easy corollary. Indeed let $N\triangleleft G$ be an open normal subgroup such that the restriction of the quotient map $q\colon G\to G/N$ to $H$ is injective. Then the homomorphism $H^{\bullet}(G/N,\mathbb{Z}/p\mathbb{Z})\to H^{\bullet}(H,\mathbb{Z}/p\mathbb{Z})$ induced by the composition of the inclusion map $H\to G$ and $q$ is finite by the above. Since this homomorphism factors through $H^{\bullet}(G,\mathbb{Z}/p\mathbb{Z})\to H^{\bullet}(H,\mathbb{Z}/p\mathbb{Z})$, the latter is also finite. ∎
@@
Let $G$ be a pro-$p$ group and let $H\subseteq G$ be a subgroup of order $p$. Then the homomorphism $H^{n}(G,\mathbb{Z}/p\mathbb{Z})\to H^{n}(H,\mathbb{Z}/p\mathbb{Z})$ is non-zero for infinitely many $n$.
@@
Assume that the claim is false and there is a natural number $d$ such that the image of $H^{n}(G,\mathbb{Z}/p\mathbb{Z})\to H^{n}(H,\mathbb{Z}/p\mathbb{Z})$ is zero for $n\geq d$. By Theorem 7.2, we can find a finite subset $S\subset H^{\bullet}(H,\mathbb{Z}/p)$ of homogeneous elements which generate $H^{\bullet}(H,\mathbb{Z}/p\mathbb{Z})$ as a $H^{\bullet}(G,\mathbb{Z}/p\mathbb{Z})$-module. Let $d^{\prime}$ be the maximal degree of the elements of $S$. Then $H^{n}(H,\mathbb{Z}/p\mathbb{Z})=0$ for every $n\geq d+d^{\prime}$. But, since $H$ has order $p$, $H^{n}(H,\mathbb{Z}/p\mathbb{Z})\neq 0$ for every $n$ which is a contradiction. ∎
@@
Let $G$ be a cohomologically quasi-Boolean pro-$2$ group. Then the quotient $H^{\bullet}(G,\mathbb{Z}/2\mathbb{Z})/\mathcal{N}(H^{\bullet}(G,\mathbb{Z}/2\mathbb{Z}))$ is a graded Boolean algebra.
@@
Let $H^{\bullet}(G,\mathbb{Z}/2\mathbb{Z})$ be the connected sum of a dual algebra $D^{\bullet}$ and a graded Boolean algebra $B^{\bullet}$. Since all elements of $D^{1}$ are nilpotent while no element of $B^{\bullet}$ is, we see that $D^{1}$ is the nilradical of the ring $H^{\bullet}(G,\mathbb{Z}/2\mathbb{Z})=D^{\bullet}\sqcap B^{\bullet}$. Hence the quotient $H^{\bullet}(G,\mathbb{Z}/2\mathbb{Z})/\mathcal{N}(H^{\bullet}(G,\mathbb{Z}/2\mathbb{Z}))$ is isomorphic to $B^{\bullet}$. ∎
@@
Let $G$ be a cohomologically quasi-Boolean pro-$2$ group. Let $B$ denote, up to isomorphism, the unique Boolean ring such that the associated graded Boolean algebra $B^{\bullet}$ is isomorphic to $H^{\bullet}(G,\mathbb{Z}/2\mathbb{Z})/\mathcal{N}(H^{\bullet}(G,\mathbb{Z}/2\mathbb{Z}))$ which exists by Lemma 7.7. We say that a continuous homomorphism $k\colon G\to\mathbb{Z}/2\mathbb{Z}$ is a quasi-canonical homomorphism if the image of the associated cohomology class $k\in H^{1}(G,\mathbb{Z}/2\mathbb{Z})$ under the quotient map
$H^{\bullet}(G,\mathbb{Z}/2\mathbb{Z})\to H^{\bullet}(G,\mathbb{Z}/2\mathbb{Z})/\mathcal{N}(H^{\bullet}(G,\mathbb{Z}/2\mathbb{Z}))$